OpenStudy (anonymous):

Will fan and medal! -3, 6, -9, 12, -15, . . . Which of the following represents the general term for the sequence given?

OpenStudy (anonymous):

(-1)^n3^n (-1)^n3n (-1)^(n+1)3

OpenStudy (amoodarya):

test formulas by put n=1 , 2, 3

OpenStudy (amoodarya):

if you look at positive of them you will find soon what they are

OpenStudy (anonymous):

(-1)^1*3^1=-3 (-1)^2*3^2=9 not it (-1)^1*3(1)=-3 (-1)^2*6=6 (-1)^3*3^3=-9 <-- correct (-1)^2*3=3 not it Okay i got it i think